OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (anonymous):

(12,-1)

jimthompson5910 (jim_thompson5910):

x − 2y = 14 x + 3y = 9 x − 2y = 14 -x - 3y = -9 (x-x)+(-2y-3y)=14-9 -5y=5 y=-1 x+3y=9 x+3(-1)=9 x-3=9 x=12 So the solution is (12,-1)